OpenStudy (darkigloo):

At what rate is the circumference changing when radius=6cm? The radius is increasing at a rate of 1.5cm/sec.

OpenStudy (darkigloo):

So i know : dr/dt=1.5 r=6 C=2 pi r i guess i have to find dC/dt

OpenStudy (nincompoop):

differentiation

OpenStudy (darkigloo):

can you explain how

OpenStudy (nincompoop):

dr/dt

OpenStudy (nincompoop):

circumference depends on the radius since pi is constant

OpenStudy (nincompoop):

so how would you set up your equation?

OpenStudy (darkigloo):

im not sure...would it be \[\frac{ d }{ dt } [ C = 2 \pi r]\]

OpenStudy (nincompoop):

go on

OpenStudy (darkigloo):

\[\frac{ dC }{ dt } = 2 \pi \frac{ dr }{ dt }\] i'm pretty sure that's incorrect...

OpenStudy (nincompoop):

since r is a function of time dC/dt = 2pir dr/dt would this be okay?

OpenStudy (darkigloo):

why does the r stay if it is being differentiated?

OpenStudy (nincompoop):

implicit differentiation

OpenStudy (nincompoop):

y'y <--

OpenStudy (nincompoop):

r'(t) = ? when r(t) = 6cm

OpenStudy (campbell_st):

I would have thought the rate of change in circumference with respect to radius is a constant and you can use related rates to solve the problem \[\frac{dC}{dt} = \frac{dC}{dr} \times \frac{dr}{dt}\]

OpenStudy (nincompoop):

C' = 2pi r'

OpenStudy (campbell_st):

so \[\frac{dC}{dt} = 2\pi \times 1.5 \]

OpenStudy (campbell_st):

since if \[C = 2\pi r\] \[\frac{dC}{dr} = 2\pi\]

OpenStudy (darkigloo):

so what happened to r=6cm?

OpenStudy (campbell_st):

well the rate of change in circumference is always 3pi no matter what the radius is... but then I'm in Year 11...

OpenStudy (campbell_st):

oops should be \[\frac{dC}{dt} = 3 \pi ..cm/ second\]

OpenStudy (darkigloo):

hmm...alright. thank you.

OpenStudy (campbell_st):

@nincompoop may have something I haven't studied yet... whatever implicit differentiation is

OpenStudy (darkigloo):

ok. i do believe the answer is 3pi. 2pi(1.5)=3pi

OpenStudy (nincompoop):

laughing out loud campbell, i was gonna show that rates can be formulated even without any given data by obtaining the general formula

OpenStudy (nincompoop):

I was looking for something that might help explain a cumbersome lesson into simple parts http://www.mathscoop.com/calculus/derivatives/applications/related-rates.php

OpenStudy (campbell_st):

ok... sorry to interrupt.... I just see these sorts of questions as related rates... sorry again..

OpenStudy (nincompoop):

you've been a help in fact

OpenStudy (darkigloo):

Thank you both.

OpenStudy (nincompoop):

thanks @campbell_st I hope the link I posted will help also, @darkigloo I had a feeling you're still at a loss with the steps

OpenStudy (campbell_st):

Hope I didn't confuse you @darkigloo

OpenStudy (darkigloo):

@nincompoop yeah i'm going to take a look at the link. thanks. @campbell_st don't worry, you didn't confuse me.